Kenley101:

A woman can bicycle 66 miles in the same time as it takes her to walk 12 miles. She can ride 9 mph faster than she can walk. How fast can she walk?

jacob28:

is der choices?

Kenley101:

No

jacob28:

k

Mercury:

in general: distance = rate*time so if you divide both sides by rate, you get time = distance/rate

Mercury:

now, let's let b = her bicycling speed and w = her walking speed for bicycling, the time equation becomes time = 66/b and for walking: time = 12/w since the time for both is equal, we can set both equations equal to each other to get 66/b = 12/w

Mercury:

finally: "She can ride 9 mph faster than she can walk" so b = w + 9 if we plug this into the equation from the previous step: 66/b = 12/w becomes 66/(w+9) = 12/w can you try solving this now? you'll need to start by cross-multiplying